OpenStudy (anonymous):

how do i get the anti- derivative for f(x)=x^2

OpenStudy (agent0smith):

Increase the power by one, and divide by the new power.

OpenStudy (anonymous):

Can i have a step-by- step ... Or apply it to another example??

OpenStudy (agent0smith):

if \[\Large f(x) = x^n\] then \[\Large F(x) = \frac{ x^{n+1} }{ n+1 }\]

OpenStudy (agent0smith):

If \[\Large f(x) = 2x^3\] then \[\Large F(x) = \frac{ 2x^{3+1} }{ 3+1 }= \frac{ 2x^4 }{ 4 }\]

OpenStudy (anonymous):

thnx